ELECTRONIC SPECTROSCOPY OF THE 6p ← 6s TRANSITION IN Au-Ne

\underline{ADRIAN M. GARDNER}, RICHARD J. PLOWRIGHT, CAROLYN D. WITHERS, TIMO-THY G. WRIGHT, School of Chemistry, University of Nottingham, University Park, Nottingham, NG7 2RD, United Kingdom; MICHAEL D. MORSE and W. H. BRECKENRIDGE, Department of Chemistry, 315 South 1400 East, Room 2020, University of Utah, Salt Lake City, Utah 84112.

Electronic spectra of the Au-Ne complex have been recorded for the first time in the vicinity of the atomic Au 6p ← 6s transition.^{a} A structured spectrum was observed near the Au ^{2}P_{3/2} ← ^{2}S_{1/2} transition, however, the complex spectrum expected near the Au ^{2}P_{1/2} ← ^{2}S_{1/2} transition was not observed. This is rationalized using high level ab intio calculations.

aR. J. Plowright, A. M. Gardner, C. D. Withers, T. G. Wright, M. D. Morse and W. H. Breckenridge, J. Phys. Chem. A, 114, 3103, (2010)